# Outcome-Refining Process Supervision for Code Generation

### v1

@@ Abstract
Large Language Models have demonstrated remarkable capabilities in code generation, yet they often struggle with complex programming tasks that require deep algorithmic reasoning. While process supervision through learned reward models shows promise in guiding reasoning steps, it requires expensive training data and suffers from unreliable evaluation. We propose Outcome-Refining Process Supervision, a novel paradigm that treats outcome refinement itself as the process to be supervised. Our framework leverages concrete execution signals to ground the supervision of reasoning steps, while using tree-structured exploration to maintain multiple solution trajectories simultaneously. Experiments demonstrate that our approach enables even smaller models to achieve high success accuracy and performance metrics on competitive programming tasks, creates more reliable verification than traditional reward models without requiring training PRMs. Our approach achieves significant improvements across 5 models and 3 datasets: an average of 26.9% increase in correctness and 42.2% in efficiency. The results suggest that providing structured reasoning space with concrete verification signals is crucial for solving complex programming tasks. [^1] We open-source all our code and data at https://github.com/zhuohaoyu/ORPS
@@
Outcome-Refining Process Supervision for Code Generation
Zhuohao Yu1*, Weizheng Gu1*, Yidong Wang1, Zhengran Zeng1, Jindong Wang2, Wei Ye1$\dagger$, Shikun Zhang1 1Peking University 2Microsoft Research zyu@stu.pku.edu.cn, wye@pku.edu.cn
@@ 1 Introduction
Figure 1: Comparison of code generation paradigms.
Large Language Models (LLMs) have demonstrated remarkable capabilities in code generation tasks (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a; Guo et al., 2024). However, when implementing algorithms with multiple components or handling intricate logic flows, these models often struggle to maintain reliability and consistency (Jiang et al., 2024b; Jimenez et al., 2023). This limitation becomes particularly apparent in problems requiring deeper algorithmic insights and careful consideration of trade-offs.
As shown in Figure 1, traditional approaches to improving LLM performance have relied on outcome supervision, where models are guided solely by final output quality (Chen et al., 2021b; Chang et al., 2023). Process supervision has emerged as a promising alternative, guiding models through intermediate reasoning steps using Process Reward Models (PRMs) (Lightman et al., 2023; Wang et al., 2024c). While effective for mathematical reasoning, its application to code generation faces unique challenges: PRMs require extensive human-annotated data for training (Luo et al., 2024; Wang et al., 2024b) and often suffer from hallucination when evaluating complex reasoning steps (Wang et al., 2023b; Chen et al., 2024a). Recent studies show that LLMs cannot reliably self-correct (Huang et al., 2023) or self-validate without external verification (Stechly et al., 2024).
@@
Code generation presents a unique opportunity through concrete, verifiable signals. Unlike other domains where intermediate steps may be difficult to verify, code can be executed throughout development, providing objective feedback about both theoretical correctness and practical performance (Zhang et al., 2023; Shinn et al., 2024). Yet existing approaches using execution feedback (Shinn et al., 2024; Zhong et al., 2024) focus primarily on local improvements and debugging, missing opportunities for exploring fundamentally different algorithmic strategies.
@@
We propose Outcome-Refining Process Supervision, a novel paradigm that treats the reasoning on refinement of outcomes itself as the process to be supervised. This approach differs fundamentally from existing self-improvement methods that focus on iterative refinement with execution feedback. Through a tree-structured exploration space, our framework maintains multiple reasoning trajectories simultaneously, enabling models to discover and refine diverse solution strategies. This structure allows models to explore different algorithmic approaches when initial attempts prove suboptimal, rather than being confined to local improvements.
Our key insight is that execution feedback can serve as objective anchors for evaluating reasoning quality, eliminating the need for specially trained PRMs. This creates a natural synergy: execution outcomes ground the evaluation process, while the model’s inherent reasoning capabilities guide the exploration of theoretical improvements. By maintaining multiple solution paths, the framework can explore diverse algorithmic strategies while ensuring each step is verified through concrete signals. This approach differs fundamentally from traditional reward models by grounding supervision in verifiable outcomes rather than learned judgments.
Through extensive experiments, we have observed several critical insights:
  • Reasoning space over model size: Providing sufficient reasoning space is more crucial than model size for complex programming tasks - even smaller models like Qwen-7B achieve remarkably high success rates (80% Pass@1) when given room to explore and refine multiple solution strategies
  • Reliable verification: Combining execution feedback with self-critique mechanisms creates a more reliable verification system than traditional reward models, without requiring expensive training data
  • Scalable improvement: Our approach shows consistent improvements in both success rates and solution efficiency compared to existing methods, particularly on complex tasks where current approaches struggle
@@
  • We propose ORPS: a novel framework that generalizes outcome and process supervision to tackle complex code problems.
  • We are the first to demonstrate that introducing search trees significantly enhances the ability to solve complex coding problems.
  • ORPS  achieves an average Pass@1 improvement of 26.9% across three datasets and five models, while reducing running time by 42.2% on average.
@@
2 Preliminaries
2.1 Process Supervision vs Outcome Supervision
Outcome supervision, the traditional paradigm in machine learning, assesses performance solely on final outputs (Chen et al., 2021b; Brown et al., 2020). This approach has evolved through LLM-as-a-judge methods, where language models are prompted (Zheng et al., 2023; Yu et al., 2024a) or trained (Wang et al., 2023c) to evaluate model outputs (Chang et al., 2023). While providing richer feedback than simple metrics, these methods still focus exclusively on final results.
Process supervision represents a paradigm shift by evaluating the quality of intermediate reasoning steps (Lightman et al., 2023; Luo et al., 2024). This paradigm employs specially trained Process Reward Models (PRMs) - an extension of LLM-as-a-judge - to evaluate each solution step (Lightman et al., 2023; Ma et al., 2023; Luo et al., 2024; Jiang et al., 2024a). PRMs have proven particularly effective in domains requiring complex reasoning, such as mathematical problem-solving and logical deduction, where they guide search algorithms toward better solution paths (Wang et al., 2024c, b).
Formally, outcome supervision evaluates only the final output: $R_{\text{outcome}}(y)=\mathbf{1}[\text{output is correct}]$. Process supervision, in contrast, aggregates step-wise scores:
$R_{\text{process}}(s_{1},...,s_{T})=\sum_{t=1}^{T}\text{PRM}(s_{t}|s_{1:t-1}),$ (1)
where $\text{PRM}(s_{t}|s_{1:t-1})$ evaluates the quality of step $s_{t}$ given the previous steps.
However, current approaches to process supervision face significant challenges in practice. The requirement for dense human annotations to train reliable PRMs makes the approach expensive and time-consuming (Lightman et al., 2023). The generalization capability of PRMs is often limited, as reasoning patterns can vary significantly across different tasks and domains. Furthermore, when serving as judges, LLMs may produce unreliable evaluations due to hallucination (Hu et al., 2024; Li et al., 2024), particularly for complex tasks (Thakur et al., 2024). Recent studies show that LLMs cannot reliably self-correct (Huang et al., 2023) or self-validate without external verification (Stechly et al., 2024). These limitations motivate our approach of grounding process supervision in concrete, verifiable signals rather than learned judgments.
@@ 2.2 Execution-Driven Code Generation
Consider implementing an efficient sorting algorithm: a model might write code that passes all test cases but uses an inefficient $O(n^{2})$ approach. Outcome supervision would mark this as a success, missing the opportunity to guide the model toward a more optimal $O(n\log n)$ solution. Similarly, if the code fails, a sparse "fail" signal provides no insight into whether the error lies in the algorithmic approach, the implementation details, or edge case handling. These limitations of both process and outcome supervision highlight the need to rethink how to supervise the development of complex programs, where both theoretical understanding and practical implementation must evolve together.
@@ 3 Methodology
We formulate code generation as a step-by-step reasoning process, where each state on our reasoning tree contains a combination of theoretical reasoning, code implementation, execution outcomes and step-level reward. This unified framework generalizes both outcome and process supervision by treating the refinement of outcomes as the process to be supervised. The overall approach is outlined in Figure 2 and formalized in Algorithm 1.
@@
0:  Problem $x$, Model $\mathcal{M}$, Beam size $K$, Steps $T$, Candidates $N$
@@
        $\texttt{candidates}\leftarrow\mathcal{M}_{\text{reason}}(x,\texttt{chain},N)$Generate N reasoning-implementation pairs to be considered as the next step}
@@
     $\texttt{beam}_{t}\leftarrow$ Select top-$K$ paths by reward {Keep promising paths}
@@
  return  Best reasoning chain from $\texttt{beam}_{T}$
3.1 Outcome-Refining Process Supervision
Our framework introduces a novel supervision paradigm where the refinement of outcomes serves as the process to be supervised. Unlike traditional approaches that rely on linear Chain-of-Thought structures, we use beam search over tree structure where each state embodies the dual nature of code generation - theoretical understanding and practical implementation. The states evolve through a self-refining process: reasoning chains capture the changes and observations on theoretical approach and implementation strategy, while execution outcomes provide concrete signals for refinement. This interaction between reasoning and execution creates a refinement process: execution outcomes inform theoretical improvements, while enhanced reasoning guides better implementations.
The tree structure, combined with beam search, enables deeper reasoning exploration crucial for complex programming tasks. By maintaining multiple promising trajectories simultaneously, our framework allows diverse exploration of different strategies - a capability particularly important for problems requiring extensive reasoning and algorithmic insights. Through comprehensive state representation including reasoning chains, code, and execution feedback, inferior solution paths are naturally replaced when better alternatives are discovered. Our experimental results in Figure 4 demonstrate this outcome-guided exploration scales effectively with increased computational budget, which suggests the importance of providing sufficient reasoning space for complex programming tasks.
3.2 Self-Critic with Generative Rewarding
Traditional PRMs are trained in a discriminative manner, producing only numerical scores without explaining their evaluation rationale. This design fails to leverage language models’ inherent ability to reason about their judgments. Our framework takes a different approach by treating process reward generation as a reasoning task itself. Before assigning scores, the model first articulates its analysis of the reasoning chain and execution outcomes, considering factors like algorithmic complexity, code structure, and performance metrics. This generative approach allows the model to benefit from its own chain-of-thought process when making evaluations.
Our experiments  Table 4 also demonstrate that this generative reward mechanism outperforms traditional discriminative PRMs, even for explicitly trained ones. The improvement stems from the model’s ability to perform detailed analysis before making judgments, rather than directly mapping inputs to numerical scores.
3.3 Execution-Guided Process Reward Model
While traditional PRMs rely on expensive training data, we leverage concrete execution signals to ground the model’s judgment. Existing methods use PRMs that are trained on human-annotated data to learn evaluation criteria for reasoning steps. Ours eliminates this training requirement through a key insight: since PRM training essentially grounds the model’s judgment through human-defined criteria, we can achieve similar grounding by providing concrete execution signals directly to the language model. This allows execution outcomes to serve as objective anchors for evaluation while preserving the model’s inherent reasoning capabilities. The process rewards emerge from holistic analysis across multiple dimensions: the soundness and coherence of reasoning steps, the relationship between reasoning and practical implementation, and concrete execution metrics including correctness, performance, and resource usage. This grounded approach offers eliminates PRM training costs, reduces hallucination through concrete verification, and enables reliable evaluation through objective metrics.

### v2

# Reasoning Through Execution: Unifying Process and Outcome Rewards for Code Generation

@@ Abstract
Large Language Models excel at code generation yet struggle with complex programming tasks that demand sophisticated reasoning. To bridge this gap, traditional process supervision relies on learned reward models requiring costly training data and suffering from reward misalignment, while outcome supervision fails for complex tasks needing coordinated intermediate steps. We introduce Outcome Refining Process Supervision, which unifies process and outcome supervision by leveraging executable verification: a tree-structured search framework generates strategic alternatives, profiles execution metrics, and scores candidates via self-critique mechanisms that integrate runtime feedback with reasoning. Experiments across 5 models and 3 benchmarks show consistent gains, with 26.9% higher correctness and 42.2% improved code efficiency. The results demonstrate that ORPS enables LLMs to overcome local optima in code generation, suggesting a promising direction for combining verifiable outcomes with structured reasoning to tackle complex challenges. We open-source at: https://github.com/zhuohaoyu/ORPS
Code Generation, Process Supervision, Reward Models, Reasoning, Large Language Models
@@ 1 Introduction
Figure 1: Comparison of outcome and process supervision.
Large Language Models (LLMs) have revolutionized code generation through their ability to synthesize programs from natural language specifications (Brown et al., 2020; Guo et al., 2024). However, complex programming tasks requiring multi-step algorithmic reasoning—such as implementing dynamic programming solutions or optimizing parallel computation patterns—remain challenging (Jiang et al., 2024b; Jimenez et al., 2023). These limitations persist even in state-of-the-art models (OpenAI, 2023; Touvron et al., 2023a), revealing a critical gap in current supervision paradigms.
As shown in Figure 1, traditional approaches follow two main paradigms: outcome supervision, which evaluates only final outcome quality (Chen et al., 2021b), and process supervision, which guides intermediate steps using learned Process Reward Models (PRMs) with search algorithms (Lightman et al., 2023). While PRMs have shown success in mathematical reasoning (Wang et al., 2024c; Chen et al., 2024a), their application to code generation faces fundamental challenges: (1) PRMs require expensive human annotations or distillations of other models on intermediate steps to train (Wang et al., 2024b); (2) Learned rewards can suffer from hallucination (misjudging invalid steps as correct) (Huang et al., 2023; Stechly et al., 2024) and reward hacking (exploiting superficial patterns to maximize scores) (Skalse et al., 2022); (3) Code-specific PRMs are scarce, and math-focused PRMs may not suit programming’s structured logic. More fundamentally, our work questions the necessity of separately trained PRMs when LLMs’ intrinsic reasoning capabilities can be effectively guided by verifiable outcomes.
@@
Code generation offers a unique advantage through concrete, verifiable signals: code can be executed throughout development, providing objective feedback on correctness and performance (Zhang et al., 2023; Shinn et al., 2024). However, existing execution-feedback methods like Reflexion (Shinn et al., 2024), LDB (Zhong et al., 2024), Self-Repair (Olausson et al., 2023), and REx (Tang et al., 2024a) primarily focus on local code repair, missing opportunities to explore fundamentally different algorithmic strategies.
@@
We propose Outcome-Refining Process Supervision, which treats the reasoning about outcome refinement itself as the process to be supervised. Unlike repair-focused methods that incrementally fix code blocks, ORPS operates at a higher abstraction level, guiding LLMs to reason about overall solution strategies. Through tree-structured exploration, our framework maintains multiple reasoning trajectories simultaneously, enabling models to explore different algorithmic approaches when initial attempts prove suboptimal (e.g., switching from brute-force to divide-and-conquer approaches), rather than being trapped in local optima (e.g., a brute-force solution that passes the test but is inefficient).
Our key insight is that LLMs’ intrinsic reasoning and self-critique capabilities, when anchored by verifiable execution feedback, can generate high-quality process rewards, eliminating the need for specially trained PRMs. Execution outcomes and performance metrics provide objective grounding, while self-critique offers strategic guidance. This creates a powerful, inference-only feedback loop where reasoning chains and code implementations co-evolve: execution failures prompt strategic re-analysis, while refined algorithmic insights lead to better implementations.
ORPS is the first inference-only process supervision framework for code generation that systematically explores algorithmic strategies without PRM training. Experiments across 5 models and 3 benchmarks reveal:
  • Strategic Reasoning Over Model Scale: Providing sufficient reasoning space for algorithmic exploration is more crucial than model size; ORPS enables smaller models to achieve high success rates, often outperforming larger counterparts using conventional methods.
  • Effective PRM Elimination: Combining execution feedback with self-critique creates superior verification mechanism compared to learned PRMs, eliminating costly training while leveraging LLMs’ inherent capabilities.
  • Algorithmic Improvement Beyond Local Optima: ORPS consistently improves correctness and solution efficiency by exploring superior algorithmic strategies rather than just performing local fixes.
@@
  • We propose ORPS: a novel inference-only framework that unifies outcome and process supervision by guiding LLMs to reason about solution strategies at a higher abstraction level, moving beyond local code repair.
  • We demonstrate that combining verifiable execution feedback with self-critique effectively eliminates the need for specially trained PRMs, outperforming them while reducing overhead.
  • ORPS achieves substantial improvements across diverse benchmarks: an average Pass@1 improvement of 26.9% across datasets and models, while reducing running time by 42.2% on average.
@@
2 Related Work
2.1 Outcome Supervision vs Process Supervision
Outcome Supervision in language models traditionally evaluates and optimizes LLM outputs through three primary paradigms: (1) Open-ended generation tasks (e.g. instruction following) assessed via text similarity metrics with reference answers (Zhang et al., 2019) or LLM-as-a-judge scoring (Zheng et al., 2023; Wang et al., 2023c; Yu et al., 2024a); (2) Constrained-output tasks (multiple-choice, math problems) judged by exact answer matching (Brown et al., 2020; Hendrycks et al., 2020) of answer keys or solutions; and (3) Code generation where correctness depends on test case execution with generated programs (Chen et al., 2021b; Austin et al., 2021). These outcomes are then used to curate the data (e.g., selecting higher-quality conversations or code solutions) for further training of the model (Liu et al., 2023b). Such approaches share a critical limitation: they ignore the reasoning process that produced the output (Uesato et al., 2022). This proves particularly problematic for complex tasks where optimal solutions require coordinated intermediate steps (Lightman et al., 2023).
Process Supervision addresses this gap by optimizing intermediate reasoning trajectories through specially trained Process Reward Models (PRMs) that score each intermediate step (Uesato et al., 2022; Lightman et al., 2023). PRMs have proven particularly effective in domains requiring complex reasoning, such as math problems, where they guide search or sampling algorithms toward better reasoning trajectories and solutions (Wang et al., 2024c; Chen et al., 2024a; Wang et al., 2024b). While outcome supervision gives $R_{\text{outcome}}(y)=\mathbf{1}[\text{correct}]$, process supervision gives step-wise reward signals $R_{\text{process}}=\sum_{t=1}^{T}\text{PRM}(s_{t}|s_{1:t-1})$.
This approach has been predominantly used in math reasoning tasks(Luo et al., 2024; Jiang et al., 2024a), but with limitations: 1) The requirement for dense human annotations to train reliable PRMs makes the approach expensive and time-consuming (Lightman et al., 2023). 2) The generalization capability of PRMs is often limited, as reasoning patterns can vary significantly across different tasks and domains. 3) When serving as judges, LLMs may produce unreliable evaluations due to hallucination (Hu et al., 2024; Li et al., 2024), particularly for complex tasks (Thakur et al., 2024). Recent studies show that LLMs cannot reliably self-correct (Huang et al., 2023) or self-validate without external verification (Stechly et al., 2024). However, emerging work suggests that specially trained PRMs may be unnecessary when LLMs’ intrinsic reasoning capabilities are effectively guided by verifiable outcomes (Chakraborty et al., 2025). These limitations motivate our approach of grounding process supervision in concrete, verifiable signals rather than learned judgments.
@@ 2.2 Execution-Driven Code Generation
Self-Repair (Olausson et al., 2023) demonstrated that reasoning quality and self-repair capabilities correlate with code generation outcomes, establishing the importance of structured reasoning through repair trees. REx (Tang et al., 2024a) further explored code repair as an exploration-exploitation tradeoff. However, these execution-guided approaches primarily focus on local code repair—debugging specific functions or fixing syntax errors—missing opportunities to explore fundamentally different strategies.
@@
Consider implementing an efficient sorting algorithm: a model might write code that passes all test cases but uses an $O(n^{2})$ approach. Existing outcome supervision methods would mark this as a success, missing the opportunity to guide the model toward an optimal $O(n\log n)$ solution. Similarly, if the code fails, a sparse ”fail” signal provides no insight into whether the error lies in the algorithmic approach, the implementation details, or edge case handling. These limitations highlight the need to rethink how to supervise the development of complex programs, where both theoretical understanding and practical implementation must evolve together—operating at a higher abstraction level than typical repair-focused methods.
@@ 3 Methodology
We propose a framework that unifies process supervision with outcome supervision by combing reasoning, code implementation, and execution verification into a single tree-structured search process.
When tackling coding tasks, particularly complex ones, it is challenging for a model to generate a fully correct solution on the first attempt. Instead, LLMs often produce imperfect yet heuristically valuable code, requiring iterative self-correction to eventually arrive at a correct implementation. Each iteration of code refinement can be considered a step in the problem-solving process.
As formalized in Algorithm 1 and illustrated in Figure 2, at step $t$, the node in a given search beam represents a state $s_{t}=(\mathcal{R}_{t},C_{t},F_{t},\omega_{t},K_{t},\rho_{t})$, where $\mathcal{R}_{t}$ denotes the current reasoning chain, $C_{t}$ the code implementation, $F_{t}$ the execution feedback, $\omega_{t}$ the outcome reward score, $K_{t}$ the self-critic reasoning and $\rho_{t}$ the process reward score.
The search progresses through three phases :
  1. 1.
    Candidate Generation. The LLM is leveraged to do reasoning on how to refine or alternative strategies and then attempt to generate corresponding code implementations.
  2. 2.
    Execution & Profiling. Each candidate code is executed and profiled on unit tests, to measure key performance metrics, such as correctness, efficiency, and code quality.
  3. 3.
    Self-Critic & Process Rewarding. The LLM is prompted to generate a texual self-critic considering reasoning chain and execution metrics, then gives a numerical process reward score which is used to guide searching.
Unlike linear CoT approaches that commit to a single trajectory, this tree structure enables parallel exploration of divergent strategies—for instance, maintaining both greedy and dynamic programming approaches for optimization problems until empirical feedback identifies the superior solution.
@@
Input: Problem $x$, Unit Tests $U$, Model $\mathcal{M}$, Beam size $K$, Steps $T$,  Candidates $N$, Process Reward Weight $\alpha$, Outcome Reward Weight  $\beta$
@@
      $\texttt{candidates}\leftarrow\mathcal{M}_{\text{reason}}(x,\texttt{chain},N)$Generate N pairs}
@@
[…]
        feedback, outcome_rew $\leftarrow$ execute_and_profile(code, $U$)
@@
    $\texttt{beam}_{t}\leftarrow$ Select Top-$K$ paths with highest step_score
@@
Return: Best reasoning chain from $\texttt{beam}_{T}$
3.1 Candidate Generation
The candidate generation phase expands the search tree by producing diverse refinements. For each node $s_{t-1}=(\mathcal{R}_{t-1},C_{t-1},F_{t-1},\omega_{t-1},K_{t-1},\rho_{t-1})$, the LLM generates $N$ successor candidates:
$\{(r_{t}^{(j)},c_{t}^{(j)})\}_{j=1}^{N}=\mathcal{M}(\mathcal{R}_{t-1},C_{t-1},F_{t-1}),$
where $r_{t}^{(j)}$ represents incremental reasoning updates (e.g., ‘Adjust loop termination to prevent off-by-one errors’) or strategic pivots (e.g., ‘Replace recursion with iteration to avoid stack overflow’). This dual-generation mechanism ensures algorithmic diversity—maintaining competing approaches until empirical feedback resolves ambiguities.
3.2 Execution & Profiling Outcome Rewards
We execute the model-generated code $\{c_{t}^{(j)}\}_{j=1}^{N}$ on a set of predefined Unit tests $U$ and execute them to obtain results. $U$ may be either generated by the LLM itself based on the problem or provided by the dataset. We will discuss the results under different scenarios in subsection 4.1.
The outcome reward $\omega_{t}^{(j)}$ combines dynamic analysis (Ball, 1999) and static analysis (Nielson et al., 1999) metrics into a unified score, grounding solution quality in both runtime behavior and structural properties. We compute:
$\omega_{t}^{(j)}=\sum_{k=1}^{M}\beta_{k}\cdot\text{normalize}(m_{k}^{(j)}),$
We have correctness, execution time, CPU instruction count, page faults as dynamic analysis metrics and code length, AST node count, cyclomatic complexity and cognitive complexity as static analysis metrics. Due to space limitations, we put detailed introduction to each metric in Table 6.
This framework rewards solutions that balance functional correctness with comprehensive aspects—a brute-force implementation passing all tests would score highly in correctness but poorly in complexity metrics, incentivizing refinement toward optimal algorithms.
3.3 Self-Critic & Process Rewarding
The same model $\mathcal{M}$ serves dual roles: generating reasoning along with solution candidates, and judging their viability. After executing $c_{t}^{(j)}$ to obtain $F_{t}^{(j)}$, the model produces a texual critique $k_{t}^{(j)}$ and a numerical process reward $\rho_{t}^{(j)}$:
$(k_{t}^{(j)},\rho_{t}^{(j)})=\mathcal{M}(\mathcal{R}_{t-1}\oplus r_{t}^{(j)},C_{t-1}\oplus c_{t}^{(j)},F_{t}^{(j)}).$
The process reward grounds subjective evaluation of the reasoning process by combining objective execution metrics. This hybrid scoring prevents reward hacking—a model cannot inflate rewards without corresponding improvements in verifiable execution outcomes. Candidates with high $\rho$ but low $\omega$ signal outcome-process mismatches, which is a critical failure mode in conventional process supervision.
3.4 Unifying Process and Outcome Supervision
The beam search mechanism proceeds with Top-K successor states using a weighted step score: $q_{t}=\alpha\rho_{t}+\beta\omega_{t}$. Where $\alpha+\beta=1$ governs the trade-off between theoretical soundness ($\rho_{t}$) and empirical effectiveness ($\omega_{t}$). The framework balances functional correctness while preserving promising reasoning trajectories.
This formulation generalizes conventional supervision paradigms. When $\beta=0$, the framework reduces to pure process supervision akin to mathematical reasoning approaches that prioritize stepwise correctness over answers during reasoning. Conversely, $\alpha=0$ recovers outcome supervision’s focus on final code quality, similar to Best-of-N sampling but on a tree. Our unified perspective reveals these as endpoints on a continuum of supervision strategies, with the proposed $\alpha/\beta$ balance enabling simultaneous optimization of reasoning and outcome solution quality.
The synergy emerges through bidirectional feedback: execution outcomes ground reasoning process by identifying discrepancies between intended and actual behavior (e.g., test failures revealing flawed base cases in recursive algorithms), while process rewards guide exploration toward algorithmically superior implementations (e.g., recognizing that memoization could transform an $O(2^{n})$ brute-force solution into an $O(n)$ dynamic programming approach).
4 Experiments
LBPP HumanEval MBPP
(Matton et al., 2024) (Chen et al., 2021b) (Austin et al., 2021)
# Test Problems 162 164 257
# Unit Tests 5.1 6.5 3.0
Solution Length§ 627 / 3039 169 / 622 130 / 589
Contamination New Dataset 18.9% 20.8%
Task Type Algorithms Func. Completion Basic Prog.
4.1 Experimental Setup
Datasets. We evaluate on 3 programming benchmarks as shown in Table 1. LBPP is a recent complex programming dataset manually curated by human experts with competitive programming experience. HumanEval and MBPP are popular code generation benchmarks but could be trivial for current LLMs (Matton et al., 2024). Moreover, a significant proportion of the data is leaked in multiple pre-training corpora (Riddell et al., 2024). To ensure reproducibility, we report our detailed hyperparameters in  Appendix A, we also open-source all our code and scripts.
Unit Tests. Our framework utilizes unit tests to verify and profile code solutions. We use LLM to generate these unit tests, which are then employed in computing the outcome reward $\omega_{t}^{(j)}$. However, self-generated unit tests do not always effectively assess code quality. Consequently, some prior works (Zhong et al., 2024) directly utilize dataset-provided unit tests. To facilitate a fair comparison with related approaches, we also consider using unit tests provided by datasets. In Table 2, methods using unit tests from datasets are denoted by (w/ T).
Baselines. We compare several strong baselines for code generation. For outcome supervision, Reflexion (Shinn et al., 2024) is a recent self-improvement approach that utilizes execution results to refine generated code. LDB (Zhong et al., 2024) extends this by incorporating debugger outputs, and intermediate variable values for iterative solution refinement. For test-time scaling, we implement Best-of-N sampling, which generates multiple solutions and selects the best one based on test outcomes. Since no existing process supervision methods have been designed specifically for code generation, we adapt a similar approach from mathematical reasoning (Luo et al., 2024) in comparison, which we include in our ablation studies.